# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v3

@@ Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights. To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing in the nonlinear setting without assumptions on the boundedness of the weight matrices. Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures. Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
@@ 1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning [25, 39], personalised recommendation systems [7, 26], social network analysis [10], and 3D mesh classification [37]. Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
@@
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth. This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing [19, 23].
@@
Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants [24, 4, 2, 21, 8, 45, 30]. At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
@@
Therefore, these metrics are often considered as sufficient but not necessary evidences for oversmoothing [33]. Despite this, there is a considerable body of literature using these unreliable metrics as their evidences for non-occurrence of oversmoothing in GNNs [47, 6, 35, 44, 20, 22, 34, 9, 36, 43, 29, 42].
@@
As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs [44, 35, 34, 45, 29, 42], where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero. Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
@@
Furthermore, we note that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix. This double implication presents an issue: in realistic settings with a large but not excessively large number of layers, we may observe the decay of the oversmoothing metric by, say, two orders of magnitude while still being far from zero. In such cases, it is unclear whether the features are aligning with the dominant eigenspace, simply decreasing in magnitude, or exhibiting neither of the two behaviours. As a result, these types of metrics provide little to no explanation for the degradation of GNN performance.
@@
As an alternative to address these shortcomings, we advocate for the use of a continuous approximation of the rank of the network’s feature representations to measure oversmoothing. Our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
@@
  • We propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights. To the best of our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
  • We provide extensive numerical evidence that continuous rank relaxation functions better capture oversmoothing than commonly used Dirichlet-like metrics.
@@ 2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees: $A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2}$, where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the node feature update of Graph Convolutional Network (GCN) [17] follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
@@ 2.2 Graph Attention Network
Graph Attention Networks (GATs) [41, 3] perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$. The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
@@ 3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN: $X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}$. Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector. Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$ [40].
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
@@ 3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus typically quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features [4, 33]
@@
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and [4] have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero [45]
@@
where $\mathcal{P}=uu^{\top}$ is the projection matrix on the space spanned by the dominant eigenvector $u$ of the matrices $A^{(l)}$. In particular, this metric can be used only if the dominant eigenvector $u$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$.
@@ 3.2 A Unifying Perspective Based on the Eigenvectors of Nonlinear Activations
In the following discussion, we present a unified and more general perspective of the necessary conditions to have oversmoothing in the sense of the classical metrics, based on the eigenvectors of a nonlinear activation function. In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix A.
@@
With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metric $E_{\mathrm{Proj}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$. Specifically, Theorem 3.2 shows that this holds provided all matrices $A^{(l)}$ share a common dominant eigenvector $u$, which is also an eigenvector of $\sigma$.
This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature. For example, (a) the proof by [4] applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by [45] holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
@@
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A_{l}$ for any $l$ and is also an eigenvector of the activation function $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
@@
Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$. Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by [4] as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$. Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$. Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$. If $\sigma=\otimes\psi$ is a nonlinear activation function acting entry-wise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$. Therefore, Theorem 3.2 implies that if the weights are sufficiently small, the features align independently of the activation function used. This is consistent with the results in [45]. However, we note that the bounds on the weights required by Theorem 3.2 and those in [45] on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant. Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs. In addition to offering a different and unifying theoretical perspective on the results in [4, 45], we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.

### v4

# Are We Measuring Oversmoothing in Graph Neural Networks Correctly?

@@ Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. We argue that these metrics have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks, while typical GNNs show a performance drop already with as few as 10 layers. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide extensive numerical evaluation across diverse graph architectures and datasets to show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that drops in the rank align closely with performance degradation, even in scenarios where energy metrics remain unchanged. Along with the experimental evaluation, we provide theoretical support for this approach, clarifying why Dirichlet-like measures may fail to capture performance drop and proving that the numerical rank of feature representations collapses to one for a broad family of GNN architectures.
@@ 1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning (Tian et al., 2022; Peng et al., 2023), personalised recommendation systems Peng et al. (2022); Damianou et al. (2024), social network analysis Fan et al. (2019), and 3D mesh classification Shi and Rajkumar (2020). Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
@@
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth. This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing Li et al. (2018); Nt and Maehara (2019); Wu et al. (2022); Rusch et al. (2023a); Zhao et al. (2024); Arnaiz-Rodriguez and Errica (2025).
@@
In recent years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants Oono and Suzuki (2019); Cai and Wang (2020); Bodnar et al. (2022); Nguyen et al. (2022); Di Giovanni et al. (2023); Wu et al. (2023); Roth and Liebig (2023). At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
@@
While these metrics are often considered as sufficient but not necessary evidence for oversmoothing Rusch et al. (2023a), there is a considerable body of literature using these unreliable metrics as their evidence for non-occurrence of oversmoothing in GNNs Zhou et al. (2021); Chen et al. (2022); Rusch et al. (2022); Wang et al. (2022); Maskey et al. (2023); Nguyen et al. (2023); Rusch et al. (2023b); Epping et al. (2024); Scholkemper et al. (2024); Wang and Cho (2024); Roth (2024); Wang et al. (2025); Arnaiz-Rodriguez and Errica (2025).
@@
However, as we show in Section˜6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. The vast majority of empirical studies in the literature that observe the decay of Dirichlet-like energy metrics are conducted over the layers of very deep but untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs Wang et al. (2022); Rusch et al. (2022; 2023b); Wu et al. (2023); Roth (2024); Wang et al. (2025), where the decay of the metrics is only driven by the small weight initialization. Instead, we show that when GNNs of different depths are trained with proper weight initialization, these metrics do not correlate with the model’s performance degradation.
@@
Furthermore, we argue that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix. This double implication presents an issue: in realistic settings with a large but not excessively large number of layers, we may observe the decay of the oversmoothing metric by, say, two orders of magnitude while still being far from zero. In such cases, it is unclear whether the features are aligning with the dominant eigenspace, simply decreasing in magnitude, or exhibiting neither of the two behaviours. As a result, these types of metrics provide little to no explanation for the degradation of GNN performance.
@@
As an alternative to address these shortcomings, we advocate for the use of a continuous approximation of the rank of the network’s feature representations to measure oversmoothing. Collapse of the rank of the feature representations was already considered as a cause of oversmoothing, e.g. in Guo et al. (2023b); Roth and Liebig (2023), however rank rank-based measures were never explicitly studied, i.e. they were never compared with other oversmoothing measures and their capacity in measuring oversmoothing was never quantified, theoretically and/or numerically. Our work fills this gap. Indeed, our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
@@
  • We notice that the rank can be a better metric for quantifying oversmoothing, and thereby we redefine oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We provide extensive numerical evidence that continuous rank relaxation functions provide a much more compelling measure of oversmoothing than commonly used Dirichlet-like metrics
Additionally, we investigate theoretically the causes of decay of the numerical rank. In particular, we show that both the aggregation matrices and the nonlinear activation functions can contribute to the decay. Our theoretical study is restricted to linear GNNs and nonlinear and non-negative GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function. For these models, we prove that the numerical rank of the features converges exactly to one. Such results provide theoretical support to our empirical evidence that oversmoothing may occur independently of the weights’ magnitude and align with our perspective on oversmoothing from the point of view of nonlinear activation functions.
@@ 2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees: $A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2}$, where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the node feature update of Graph Convolutional Network (GCN) Kipf and Welling (2016) follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
@@ 2.2 Graph Attention Network
Graph Attention Networks (GATs) Veličković et al. (2017); Brody et al. (2021) perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$. The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
@@ 3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN: $X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}$. Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector. Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$, see e.g. tudiscoComplexPowerNonnegative2015.
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ does not diverge in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
@@ 3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus typically quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features Cai and Wang (2020); Rusch et al. (2023a)
@@
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and Cai and Wang (2020) have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero Wu et al. (2023)
@@
where $\mathcal{P}=uu^{\top}$ is the projection matrix on the space spanned by the dominant eigenvector $u$ of the matrices $A^{(l)}$.
Note that both these metrics, $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, can be used only if the dominant eigenvector of $A^{(l)}$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$. Moreover, both these metrics essentially measure the deviation of the feature representations from the dominant eigenspace of the aggregation matrices $A^{(l)}$. So we expect them to perform very similarly in capturing oversmoothing. In particular, it is not difficult to show that they are equivalent metrics from a mathematical point of view, i.e. there exist constants $C_{1},C_{2}>0$ such that $C_{1}E_{\mathrm{Dir}}(X)\leq E_{\mathrm{Proj}}(X)\leq C_{2}E_{\mathrm{Dir}}(X)$, see Lemma˜A.1.
@@ 3.2 A Unifying Perspective Based on the Eigenvectors of Nonlinear Activations
We present here a unifying and more general perspective of the necessary conditions to have oversmoothing in the sense that $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ decay to zero, based on the concept of eigenvectors for a nonlinear activation function. In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix˜A.
@@
With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metrics $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$. Specifically, Theorem 3.2 shows that this holds provided all matrices $A^{(l)}$ share a common dominant eigenvector $u$, which is also an eigenvector of $\sigma$.
@@
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A^{(l)}$ for any $l$ and also an eigenvector of the activation $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
@@
[…]